OpenStudy (anonymous):

Determine the vector equation of each line with the same x-intercept as [x, y, z] = [3, 0, 0] + t[4, -4, 1] and the same z-intercept as [x, y, z] = [6, 3, -1] + t[3, 6, -2] The answer is: [x, y, z] = [0, 0, -3] + t[-3, 0, -3] Please help! Thanks :)

OpenStudy (anonymous):

Its corrected now

OpenStudy (ajprincess):

x=3+4t y=o+4t z=0+t For the x-intercept of L1 y=0, z=0. This happens when t=0. hence x=3. The x-intercept of is: .(3,0,0). x=6+3t y=3+6t z=-1-2t For the z-intercept of L2 x=0, y=0 0=6+3t -6=3t t=-2

OpenStudy (ajprincess):

http://www.sosmath.com/CBB/viewtopic.php?t=48215 need to go nw. that link will b of some use.

OpenStudy (anonymous):

Thanks anyway :) Bye!